OpenStudy (anonymous):

PLEASE HELP!!!! Let u = <7, -3>, v = <-9, 5>. Find 4u - 3v.

OpenStudy (jacob902):

heres an example

OpenStudy (jacob902):

(4y+2x-5)dx+(6y+4x-1)dy=0, y(-1)=2 i tried solving by seperation of variables but i didnt get newhere with that

OpenStudy (jacob902):

You can do a linear transformation of the variables to get rid of the "-5" and "-1". If 4y + 2x - 5 = 0 and 6y + 4x - 1 = 0 then y = 9/2, x = -13/2 So let x = u - 13/2 y = v + 9/2 du = dx dv = dy (4v + 2u) du = (6v + 4u) dv dv/du = (2v + u)/(3v + 2u) That is a homogeneous equation so substitute v = wu dv/du = w + u dw/du w + u dw/du = (2w + 1)/(3w + 2) Now you can separate the variables and integrate. EDIT: Duh!!!!! <face-palm> I just saw your other question You don't need to do all that stuff, it's an exact equation!!! Integral is 4xy + x^2 - 5x + 3y^2 - y = c y(-1) = 2 -8 + 1 + 5 + 12 - 2 = c c = 8

OpenStudy (anonymous):

honestly none of this makes sense to me @Jacob902